### v1

@@ IV Imperfection of the system
V Comprehensive analysis for parameter estimation
In this section we provide analysis on the practical position of maximum sensitivity, further details on amplitude estimation, and metrological gain vs measurement repetitions.
In part I, we see that ideally maximum FI is achieved at $\beta\rightarrow 0$. In the presence of imperfections, this is not the case. To exemplify this let us first compute the FI of SCS for phase estimation. The probability of overlap is given by
@@ V Comprehensive analysis for parameter estimation
$=$ $\frac{1}{4(1+e^{-\frac{\alpha^{2}}{2}})}|1+2e^{-\frac{\alpha^{2}}{2}}+e^{-\alpha^{2}(1-e^{i\theta})}|^{2}.$
@@
The supplemental analysis of amplitude estimation is shown in Fig. 8. Similar to phase estimation, it is essential to gain insights into the mean sensitivity above maximum FI of CS and the $\gamma$ offset at which the FI is optimal. With the increase of photon number, the average FI increases linearly while the optimal amplitude sensor point is moving to 0 as shown in Fig. 8(a). The optimal amplitude sensor point can also be shifted to 0 by adding a displacement to the initial state in the GRAPE pulse, which is crucial for direct amplitude sensing. Furthermore, to estimate the amplitude estimation error directly via $\Delta\gamma=\Delta p/|\partial p/\partial\gamma|$, we present a benchmark in Fig. 3(c). The amplitude estimation error for the CS remains constant even with an increase in the average photon number. In contrast, the estimation error using SCS exhibits an almost linear relationship with the average photon number and shows an optimal gain of 9.3(5)dB compared to the CS.
Furthermore, we also perform amplitude estimation with unbalanced SCS for optimization. We note two distinct scenarios: one with a fixed average photon number of 1 and another with a fixed $\alpha=1.6$, shown in Fig. 9(a) and Fig. 9(b), respectively. The former shows the less weight on the $\alpha$ component of SCS yields a higher sensitivity. In the latter, the optimal sensitivity with limited $\alpha$ is achieved by utilizing balanced components.
@@
Figure 10: Metrological gain vs measurement repetitions. The gain of SCS over CS for: (a) phase estimation and (b) amplitude estimation for different measurement repetitions.
@@
Finally, an important factor to consider for sensing applications is the measurement repetitions, which also directly translates to the total duration of the protocol. The highest metrological gain for phase and amplitude estimation we report in the main text was done with $1000$ measurement repetitions. The respective gain vs measurement repetitions are plotted in Fig. 10. Notably, the gain for both phase and amplitude estimation stabilizes at $\sim 600$ measurement repetitions. It shows that beyond this point, the single shot measurements is enough to provide meaningful statistics.

### v2

@@ Appendix D Imperfections in the system
The dashed curves in Fig. 2(d) represent simulations that account for all imperfections. To investigate the individual impact of each imperfection on the maximum FI, we perform simulations in which each imperfection is considered independently. The solid curves in Fig. 7 illustrate the theoretical QFI (pink) and the maximum FI as a function of the average photon number $\bar{n}$ for several cases using SCS: green (no decoherence and no self-Kerr), orange (cavity $T_{c,1}$), gray (self-Kerr), red (qubit $T_{q,1}$ and $T_{q,2}$), and blue (all imperfections). Notably, the green curve does not closely saturate the QFI. This can be attributed to imperfect state preparation via GRAPE pulses, which achieves a fidelity of approximately $0.998$. As expected, simulations using ideal initial SCS in place of the one generated via GRAPE pulse (not shown) confirms that the outcome saturates the QFI. Cavity decay (solid orange curve) results in a minor reduction in the FI due to the cavity’s long lifetime ($\approx 1$ ms). The self-Kerr effect, an unwanted term in the Hamiltonian that distorts the state in phase space during the phase-encoding process, causes a slightly larger decrease in the FI (solid gray curve). The most significant degradation arises from the qubit decoherence (solid red curve). During the dynamics, numerical pulses entangle the qubit and cavity states to facilitate operations on the cavity state, but at the same time making it vulnerable to qubit energy decay and dephasing. Finally, the solid blue curve incorporates all imperfections. For comparison, we also include results for squeezed vacuum states: dashed pink (QFI) and dashed blue (all imperfections). Although squeezed vacuum states theoretically provide higher QFI, they are more susceptible to decoherence experimentally, with the FI (dashed blue curve) losing its quadratic dependence and appearing almost linear under the realistic hardware imperfections considered here.
Figure 9: Amplitude estimation. (a) FI for amplitude estimation using CS (solid purple), SCS-v for amplitude in the vertical direction (solid blue), and SCS-h for amplitude in the horizontal direction (solid grey). The corresponding dashed curves indicate simulation. (b) The maximum FI for SCS-v (blue) and CS (purple) vs photon number $\overline{n}$. Dashed curves are simulation and dotted curves are fitting: $4.9(2)\overline{n}^{1.06(2)}+2.2(1)$ for SCS-v and $0.411(1)\overline{n}^{0.00(3)}$ + 2.165(1) for CS. The solid pink curve is the ideal QFI of SCS-v.
Figure 10: Performance of quantum-enhanced amplitude precision. (a) The average FI of SCS for which it is above the max FI of CS and $\gamma$ offset at which FI is maximum vs $n$. The dotted line is lineaer fit of $3.2(2)\overline{n}^{1.0(3)}+2.2(1)$. (b) The same protocol as in Fig. 4(c) in main text with initial states prepared with an offset. (c) The amplitude precision vs $n$ for both SCS (blue circles) and CS (purple diamonds). The dotted lines are linear fit of $\log(\Delta\gamma)=a\log(\overline{n})+b$ with $a$ and $b$ as free parameters. The fitted parameters are $a,b=-0.460(3),-1.311(2)$ for SCS and $a,b=-0.0105(2),-0.480(1)$ for CS.
In Appendix A, we see that ideally maximum FI is achieved at $\beta\rightarrow 0$. In the presence of imperfections, this is not the case. To exemplify this let us first compute the FI of SCS for phase estimation. The probability of overlap is given by
@@
$=$ $\frac{1}{4(1+e^{-\frac{\alpha^{2}}{2}})^{2}}|1+2e^{-\frac{\alpha^{2}}{2}}+e^{-\alpha^{2}(1-e^{i\theta})}|^{2}.$
@@
Appendix E Amplitude estimation using the superposition of coherent states
To illustrate the versatility of our protocol, we demonstrate quantum-enhanced precision for amplitude estimation by simply changing the process unitary of Fig. 2(a) in the main text to a displacement in the imaginary direction $\hat{U}(\gamma)=\exp(i\gamma(\hat{a}^{\dagger}+\hat{a}))$, where $\gamma$ is the displacement amplitude. The process generator is now $\hat{x}=(\hat{a}^{\dagger}+\hat{a})$, making states with high variance $(\Delta\hat{x})^{2}$ useful for amplitude estimation. Here, SCS offer high variance $(\Delta\hat{x})^{2}=2\overline{n}+1$ compared to CS, which have a constant variance $(\Delta\hat{x})^{2}=1$. The illustration is similar to that of phase estimation of Fig. 2(b) in the main text, except the SCS are now displaced in the imaginary direction instead of rotated. SCS with higher $\alpha$ have more rapidly oscillating fringes such that the overlap with the initial state reduces more sharply as $\gamma$ increases.
@@ Appendix E Amplitude estimation using the superposition of coherent states
Figure 11: Metrological gain vs measurement repetitions. The precision gain of SCS over CS for (a) phase and (b) amplitude estimation for different numbers of measurement repetitions.
Experimentally, we show that SCS displaced vertically in the imaginary direction (SCS-v) offer the highest FI, compared to both CS and SCS displaced horizontally in the real direction (SCS-h), as shown in Fig. 9(a). This can be explained by noting that SCS-h has a different process unitary $\hat{U}(\gamma)=\exp(i\gamma i(\hat{a}-\hat{a}^{\dagger}))$, where the process generator is now $\hat{y}=i(\hat{a}-\hat{a}^{\dagger})$, such that the variance of SCS is low, even lower than that of the CS.
Furthermore, by operating at the maximum FI, we show that the SCS-v (CS) exhibits a linear (constant) trend against $\overline{n}$, as shown in Fig. 9(b), indicating $\sqrt{\overline{n}}$ quantum enhancement in precision.
Similar to phase estimation, it is essential to gain insights into the mean sensitivity above maximum FI of CS and the $\gamma$ offset at which the FI is optimal. With the increase of photon number, the average FI increases linearly while the optimal amplitude sensor point is moving to 0 as shown in Fig. 10(a). The optimal amplitude sensor point can also be shifted to 0 by adding a displacement to the initial state in the GRAPE pulse, which is crucial for direct amplitude sensing. Furthermore, to estimate the amplitude estimation error directly via $\Delta\gamma=\Delta p/|\partial p/\partial\gamma|$, we present a benchmark in Fig. 10(c). The amplitude estimation error for the CS remains constant even with an increase in the average photon number. In contrast, the estimation error using SCS exhibits an almost linear relationship with the average photon number and shows a maximum gain of 9.3(5)dB.
@@
Finally, an important factor to consider for sensing applications is the measurement repetitions, which also directly translates to the total duration of the protocol. The highest metrological gain for phase and amplitude estimation we report in the main text was done with $1000$ measurement repetitions. The respective gain vs measurement repetitions are plotted in Fig. 11. Notably, the gain for both phase and amplitude estimation stabilizes at $\sim 600$ measurement repetitions. It shows that beyond this point, the single-shot measurements are enough to provide meaningful statistics.
Appendix F Bayesian inference
Bayesian estimation is a powerful statistical tool that updates the probability estimate for a hypothesis as more evidence or information becomes available [46]. In the context of quantum metrology, Bayesian inference allows for the optimal estimation of parameters by systematically incorporating experimental data to refine prior knowledge. Here we first briefly recap the Bayesian method and then utilize it to perform phase estimation $\theta$ with our experimental data using SCS to recover the results and trends presented in the main text. In addition, we also perform Bayesian inference to directly estimate the dispersive coupling strength $\chi$.
F.1 The Bayesian framework
The Bayesian approach hinges on Bayes’ theorem, which relates the posterior probability $P(\theta|\mathcal{D})$ of a parameter $\theta$ given data $\mathcal{D}$ to the prior probability $P(\theta)$ and the likelihood $P(\mathcal{D}|\theta)$:
$P(\theta|\mathcal{D})=\frac{P(\mathcal{D}|\theta)P(\theta)}{P(\mathcal{D})},$ (10)
where $P(\mathcal{D})$ is the evidence, ensuring the posterior is properly normalized.
In our study, the parameter of interest is the phase $\theta$, and the data $\mathcal{D}$ comprises the measurement outcomes from experiments using SCS probes.
For each experimental trial, the overlap $p_{g0}(\theta)=|\langle\psi|\psi(\beta)\rangle|^{2}$ is measured in a single-shot manner, yielding outcomes $1$ or 00 whose probabilities shall be respectively denoted henceforth as $p_{1}(\theta)=p(\theta)$ and $p_{0}(\theta)=1-p(\theta)$ for ease of notation. The likelihood function for observing $N_{1}$ number of outcome $1$ in $N$ trials is given by:
$L(\theta)=p_{1}(\theta)^{N_{1}}\,(1-p_{1}(\theta))^{N-N_{1}}.$ (11)
Assuming a uniform prior $P(\theta)$ over the interval $\theta\in[0,2.5]$, the posterior distribution simplifies to being proportional to the likelihood:
$P(\theta|\mathcal{D})\propto L(\theta).$ (12)
The Bayesian estimate, $\theta_{\text{Bayesian}}$, is obtained by maximizing the posterior distribution, and the uncertainty $\Delta\theta$ is given by the standard deviation of the posterior.
Figure 12: Comparison of the estimated phase via Bayesian inference $\theta_{\text{Bayesian}}$ with the true phase $\theta_{\text{True}}$ for various SCS amplitudes $\alpha$. Error bars represent the standard deviation of the posterior distribution. The solid lines correspond to equality.
Figure 13: Uncertainty in the estimation $\Delta\theta$ as predicted by the Cramér-Rao bound (blue) and Bayesian inference (red). Different panels correspond to various coherent amplitudes $\alpha$ of the SCS. The Bayesian estimates generally align with the Cramér-Rao bound predictions within the dynamical range, validating the efficiency of the Bayesian approach in approaching the Cramér-Rao bound.
F.2 Phase estimation
In order to get the probability $p_{1}(\theta)$, we fit the experimental data with the probability model inspired by the theoretical overlap in Eq. (9):
$p_{1}(\theta)=\frac{\left(c+e^{-{\alpha^{\prime}}^{2}\left(1-e^{-i\theta}\right)}\right)\left(c+e^{-{\alpha^{\prime}}^{2}\left(1-e^{i\theta}\right)}\right)}{4b},$ (13)
where the free parameters $b$, $c$, and $\alpha^{\prime}$ (effective amplitude of the coherent state) account for experimental imperfections. This fitted probability function, together with the single-shot experimental data $N_{1}$ at each phase $\theta$ allows the calculation of $\theta_{\text{Bayesian}}$ and the standard deviation $\Delta\theta$ from the posterior distribution given in Eq. (12).
We illustrate the alignment between the Bayesian estimates and the true phase across different $\alpha$ values in Fig. 12. As $\alpha$ increases, the precision of the estimation (within the dynamical range) improves, which is evident from the reduced error bars. However, this comes with a trade-off in the dynamical range, limiting the range of $\theta$ over which the estimation remains accurate.
Next, we compares the uncertainty $\Delta\theta$ derived from Bayesian estimation with the Cramér-Rao bound prediction via $\Delta\theta=1/\sqrt{N\,\text{FI}}$ in Fig. 13. The close agreement between the two methods, within the dynamical range, demonstrates the effectiveness of Bayesian inference in achieving near-optimal parameter estimation. The minimal uncertainty $\Delta\theta_{\text{min}}$ as a function of the coherent state amplitude $\alpha$ (or average photon number $\bar{n}$) exhibits a decreasing trend, as shown in Fig. 14, indicating enhanced precision with larger $\bar{n}$. This scaling is consistent with theoretical predictions based on the Cramér-Rao bound.
Figure 14: Minimal estimation uncertainty $\Delta\theta_{\text{min}}$ plotted against the average photon number $\bar{n}$ for various SCS. Both Bayesian estimation and Cramér-Rao bound predictions show a similar decreasing trend, confirming the scaling relationship between precision and $\bar{n}$.
F.3 Estimation of $\chi$
The phase $\theta$ is linearly related to the interaction strength $\chi$ by $\theta=\chi t/2$, assuming the time $t$ is precisely known and controlled. In this framework, the uncertainty in $\theta$ directly translates to the uncertainty $\Delta\chi$, and selecting the optimal $t$ corresponds to optimizing $\theta$ for sensitivity.
An iterative strategy can be used to refine the estimate of $\chi$. Initially, an arbitrary choice of $t^{(1)}$ is used for the first measurement round. From the resulting experimental data, a Bayesian estimate $\chi^{(1)}$ is obtained. In the subsequent iteration, the interaction time is adjusted to $t^{(2)}$, chosen such that $\chi^{(1)}t^{(2)}/2\approx\theta_{\mathrm{opt}}$, where $\theta_{\mathrm{opt}}$ is the point of maximum sensitivity (at which $\text{FI}(\theta)$ is maximum). This process repeats, with each iteration refining the estimate $\chi^{(k)}$ and updating the interaction time $t^{(k+1)}$ accordingly. The iterations continue until the desired precision in $\chi$ is achieved, leveraging Bayesian inference at each step to optimize and update the experimental parameters dynamically.
For the subsequent iteration, we can also make use of all the data accumulated in the previous rounds. As such, we estimate $\chi$ with Bayesian method utilizing the data of different $t$’s. Let us consider $N_{t}$ discrete times $\{t_{i}\}_{i=1}^{N_{t}}$. At each time $t_{i}$, we perform $N_{i}$ measurements and record the number of times the measurement yields outcome $1$ as $N_{1}[t_{i}]$, with the number of 00 detections being $N_{0}[t_{i}]=N_{i}-N_{1}[t_{i}]$. The probability of detecting the outcome $1$ at time $t_{i}$ given the interaction strength $\chi$ is
$p_{1}(t_{i}|\chi)=p_{1}\left(\theta_{i}\right)=p_{1}\left(\frac{\chi t_{i}}{2}\right),$ (14)
where $p_{1}(\theta)$ is the probability function derived via Eq. (13), as discussed earlier. The likelihood function for $\chi$ based on all the measurement outcomes is then constructed as
$\mathcal{L}(\chi)=\prod_{i=1}^{N_{t}}\left[p_{1}(t_{i}|\chi)\right]^{N_{1}[t_{i}]}\left[1-p_{1}(t_{i}|\chi)\right]^{1-N_{1}[t_{i}]}.$ (15)
Assuming a uniform prior for $\chi$ over a plausible range, the posterior probability distribution is proportional to the likelihood function:
$\mathcal{P}(\chi)\propto\mathcal{L}(\chi).$ (16)
The maximum of $\mathcal{P}(\chi)$ yields the most probable estimate $\chi_{\text{est}}$, and the standard deviation of the posterior distribution provides the uncertainty in the estimation $\Delta\chi$.
Figure 15: Convergence of the estimated interaction strength $\chi_{\text{est}}$ as a function of iteration steps in the adaptive estimation procedure. The probe states are the SCS with $\alpha=1$ and $2$, with $N=1000$ new sampled data per iteration. The rapid reduction in estimation error illustrates the efficiency of the adaptive Bayesian approach. The higher value of $\alpha$ further enhances the estimation precision, leading to faster convergence to the true value.
We illustrate the adaptive estimation process for the interaction strength $\chi$ over multiple iteration steps in Fig. 15. The estimated value rapidly converges to the independently calibrated value of $\chi$, shown as the blue line. At each iteration, the estimation leverages all accumulated data, leading to progressively improved precision. The figure highlights the robustness and efficiency of the adaptive Bayesian approach in achieving accurate and reliable estimates of $\chi$. For example, with only 4 iterations, $\Delta\chi/2\pi\approx 0.024$ MHz ($1.7\%$) for SCS with $\alpha=2$. Better precision can be obtained by using more iterations or using SCS with higher amplitude $\alpha$.